### v1

@@ 3 On $\Psi$-semismooth single-valued mappings
The following definition is closely related to the definition of semismooth operators in Banach spaces, cf. [31, Definition 3.1].
@@
    We say that $F$ is $\Psi$-ss on a subset $U\subset\Omega$ if $F$ is $\Psi$-ss at every $x\in U$.
Remark 3.2.
If $F$ is $\Psi$-ss at $\bar{x}$ then $\Psi$ is also called a Newton map for $F$ at $\bar{x}$ in [15]. If $\Psi$ is single-valued then it is also called a Newton derivative (or slanting function) for $F$ at $\bar{x}$ in the literature. We dispense with using this terminology because they are normally used at single points, namely solutions of the nonsmooth equation $F(x)=0$. However, for our purposes it is essential that the properties holds not only at one single point $\bar{x}$ but also for all $x$ belonging to some neighborhood of $\bar{x}$. In the literature one can find also several papers about the $\Psi$-ss property under different names using some additional assumptions on $\Psi$, see, e.g., [13], [22], [23].
Clearly, continuously differentiable mappings $F$ are $\nabla F$-ss.
Lemma 3.3.
Assume that $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is semismooth at $\bar{x}\in\Omega$ with respect to some mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$. Then $F$ is calm at $\bar{x}$ and $\mathrm{clm}\,F(\bar{x})\leq\mathrm{bnd}\,F(\bar{x})$.
Proof.
By definition, for every $\epsilon>0$ we can find some neighborhood $U\subset\mathrm{dom}\,\Psi(\bar{x})\cap\Omega$ of $\bar{x}$ such that for every $x\in U$ and every $A\in\Psi(x)$ there holds $\|A\|\leq\mathrm{bnd}\,\Psi(\bar{x})+\epsilon$ and $\|F(x)-F(\bar{x})-A(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|$ implying $\|F(x)-F(\bar{x})\|\leq(\|A\|+2\epsilon)\|x-\bar{x}\|$. Since $\Psi(x)\not=\emptyset$ and $\epsilon>0$ can be made arbitrarily small, the assertion follows. ∎
It is well-known that Newton maps satisfy a chain rule, see, e.g., [15, Theorem 6.14]. For the sake of completeness we state the following result, where in contrast to [15] we do not assume Lipschitz continuity of the involved mappings $F_{1},F_{2}$:
@@
By Lemma 3.3 we can find a neighborhood $\tilde{V}\subset V$ of $F_{1}(\bar{x})$ together with some real $\kappa_{2}>0$ such that $\|F_{2}(y)-F_{2}(F_{1}(\bar{x}))\|\leq\kappa_{2}\|y-F_{1}(\bar{x})\|$ and $\mathrm{bnd}\,\Psi_{2}(y)\leq\kappa_{2}$ $\forall y\in\tilde{V}$. Again by Lemma 3.3, we can find some neighborhood $\tilde{U}\subset U$ of $\bar{x}$ with $F_{1}(\tilde{U})\subset\tilde{V}$ together with $\kappa_{1}>0$ such that $\|F_{1}(x)-F_{1}(\bar{x})\|\leq\kappa_{1}\|x-\bar{x}\|$ and $\mathrm{bnd}\,\Psi_{1}(x)\leq\kappa_{1}$ for all $x\in\tilde{U}$. Next, for arbitrary $\epsilon>0$ we choose a neighborhood $V_{\epsilon}\subset\tilde{V}$ of $F_{1}(\bar{x})$ such that for each $y\in V_{\epsilon}$ and each $B\in\Psi_{2}(y)$ one has $\|F_{2}(y)-F_{2}(F_{1}(\bar{x}))-B(y-F_{1}(\bar{x}))\|\leq\epsilon\|y-F_{1}(\bar{x})\|$. Finally we choose a neighborhood $U_{\epsilon}\subset\tilde{U}$ such that $F_{1}(U_{\epsilon})\subset V_{\epsilon}$ and for every $x\in U_{\epsilon}$ and every $A\in\Psi_{1}(x)$ one has $\|F_{1}(x)-F_{1}(\bar{x})-A(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|$. Thus, for every $x\in U_{\epsilon}$, every $A\in\Psi_{1}(x)$ and every $B\in\Psi_{2}(F_{1}(x))$ we obtain
@@
Consider an open set $\Omega\subset\mathbb{R}^{n}$, two mappings $F_{i}:\Omega\to\mathbb{R}^{m}$ and two mappings $\Psi_{i}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, $i=1,2$. If $F_{i}$ is $\Psi_{i}$-ss, $i=1,2$ at $\bar{x}\in\Omega$ then $F_{1}+F_{2}$ is $(\Psi_{1}+\Psi_{2})$-ss at $\bar{x}$.
@@
Let $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, be continuous and assume that $F$ is semismooth at $\bar{x}\in\Omega$ with respect to $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$. Then $F$ is also $\rm cocl\,\Psi$-ss at $\bar{x}$.
@@
Let $\epsilon>0$ be arbitrarily fixed. Since $\Psi$ is locally bounded at $\bar{x}$, we can find some $\bar{\delta}>0$ and some $M>0$ such that $\sup\{\|A\|\,\big{|}\,A\in\bigcup_{x\in{\cal B}_{\bar{\delta}}(\bar{x})}\Psi(x)\}\leq M$. By definition we can find some $\delta\in(0,\bar{\delta})$ such that for every $x\in{\cal B}_{\delta}(\bar{x})$ and every $A\in\Psi(x)$ there holds
$\|F(x)-F(\bar{x})-A(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|.$
Now consider $x\in{\cal B}_{\delta/2}(\bar{x})\setminus\{\bar{x}\}$ and $A\in(\rm cocl\,\Psi)(x)$. Then $A$ can be written as a convex combination of elements in $\mathrm{cl}\,\Psi(x)$:
$A=\sum_{i=1}^{N}\lambda_{i}A_{i},\ \sum_{i=1}^{N}\lambda_{i}=1,\ \lambda_{i}\geq 0,\ A_{i}\in\mathrm{cl}\,\Psi(x),\ i=1,\ldots,N.$
For every matrix $A_{i}$ we can find some pair $(x_{i},B_{i})\in\mathrm{gph}\,\Psi$ with $\|A_{i}-B_{i}\|\leq\epsilon$, $\|x_{i}-x\|<\min\{\epsilon\|x-\bar{x}\|,\frac{\delta}{2}\}$ and $\|F(x)-F(x_{i})\|\leq\epsilon\|x-\bar{x}\|$. Then $x_{i}\in{\cal B}_{\delta}(\bar{x})$, $\|\bar{x}-x_{i}\|\leq(1+\epsilon)\|\bar{x}-x\|$ and
$\|F(x)-F(\bar{x})-A(x-\bar{x})\|=\|\sum_{i=1}^{N}\lambda_{i}\big{(}F(x)-F(\bar{x})-A_{i}(x-\bar{x})\big{)}\|\leq\sum_{i=1}^{N}\lambda_{i}\|F(x)-F(\bar{x})-A_{i}(x-\bar{x})\|$
$\leq\sum_{i=1}^{N}\lambda_{i}\big{(}\|F(x_{i})-F(\bar{x})-B_{i}(x_{i}-\bar{x})\|+\|F(x)-F(x_{i})\|+\|B_{i}(x_{i}-\bar{x})-A_{i}(x-\bar{x})\|\big{)}$
$\leq\sum_{i=1}^{N}\lambda_{i}\big{(}\epsilon\|x_{i}-\bar{x}\|+\epsilon\|x-\bar{x}\|+\|B_{i}\|\|x_{i}-x\|+\|B_{i}-A_{i}\|\|x-\bar{x}\|\big{)}$
$\leq\sum_{i=1}^{N}\lambda_{i}\big{(}\epsilon(1+\epsilon)\|x-\bar{x}\|+\epsilon\|x-\bar{x}\|+M\epsilon\|x-\bar{x}\|+\epsilon\|x-\bar{x}\|\big{)}=\epsilon(3+\epsilon+M)\|x-\bar{x}\|.$
Since $\epsilon>0$ can be taken arbitrary small, this inequality verifies that $F$ is $\rm cocl\,\Psi$-semismooth at $\bar{x}$. ∎
@@
The elements of $G(x)$ are called pseudosubgradients.
@@
Note that pseudosubgradients are column vectors whereas the elements of $\Psi$ are $1\times n$ matrices, i.e., row vectors.
@@
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $f:\Omega\to\mathbb{R}$ be semismooth on $\Omega$ with respect to the mapping $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{1\times n}$. Then the following statements hold true.
@@
    For every $x\in\Omega$ there holds ^$\widehat{\partial}f(x)^{T}\subset(\rm cocl\,\Psi)(x)$.
@@
    Let $x,x^{\prime}\in\Omega$ be such that the line segment $[x,x^{\prime}]$ connecting $x$ and $x^{\prime}$ is contained in $\Omega$. Then there exists $\bar{t}\in(0,1)$ and $g\in(\rm cocl\,\Psi)\big{(}x+\bar{t}(x^{\prime}-x)\big{)}$ with $f(x^{\prime})-f(x)=g(x^{\prime}-x)$.
  3. (iii)
    $f$ is locally Lipschitz on $\Omega$ and for $x\in\Omega$ there holds ${\rm lip\,}f(x)\leq\mathrm{bnd}\,\Psi(x)$.
  4. (iv)
    For all $x\in\Omega$ there holds $\partial f(x)^{T}\subset\partial_{c}f(x)^{T}\subset(\rm cocl\,\Psi)(x)$.
  5. (v)
    Let $\Omega_{\Psi}$ denote the set of all $x\in\Omega$ such that $(\rm cocl\,\Psi)(x)$ is a singleton. Then $\Omega\setminus\Omega_{\Psi}$ has Lebesgue measure zero. Moreover, the function $f$ is strictly differentiable at every $x\in\Omega_{\Psi}$ with $\{\nabla f(x)\}=(\rm cocl\,\Psi)(x)$ and the mapping $x\mapsto\nabla f(x)$ is continuous on $\Omega_{\Psi}$.
@@
ad (i): Assume on the contrary that there exist $x\in\Omega$ and $g^{T}\in\widehat{\partial}f(x)$ with $g\not\in(\rm cocl\,\Psi)(x)$ and consider the projection $\bar{g}$ of $g$ onto the convex compact set $(\rm cocl\,\Psi)(x)$, which must be different from $g$. Pick any sequence $t_{k}\downarrow 0$ and consider the points $x_{k}:=x+t_{k}(g-\bar{g})^{T}$. For sufficiently large $k$ we have $x_{k}\in\Omega$ and we can pick some $g_{k}\in\Psi(x_{k})$. Then the sequence $g_{k}$ is bounded and, by possibly passing to a subsequence, we can assume that $g_{k}$ converges to some $\hat{g}$. Clearly, $\hat{g}\in(\mathrm{cl}\,\Psi)(x)\subset(\rm cocl\,\Psi)(x)$ and therefore $(g-\bar{g})(\hat{g}-\bar{g})^{T}\leq 0$. Since $f$ is $\Psi$-ss on $\Omega$, we obtain that
$\lim_{k\to\infty}\frac{f(x_{k})-f(x)-g_{k}(x_{k}-x)}{\|x_{k}-x\|}=0.$
Further we have
$\liminf_{k\to\infty}\frac{f(x_{k})-f(x)-g(x_{k}-x)}{\|x_{k}-x\|}\geq 0$
due to $g^{T}\in\widehat{\partial}f(x)$ and consequently
$0\leq\liminf_{k\to\infty}\frac{(g_{k}-g)(x_{k}-x)}{\|x_{k}-x\|}=\frac{(\hat{g}-g)(g-\bar{g})^{T}}{\|g-\bar{g}\|}=\frac{(\hat{g}-\bar{g})(g-\bar{g})^{T}}{\|g-\bar{g}\|}-\|g-\bar{g}\|<0,$
a contradiction.
ad(ii): By Lemma 3.3, $f$ is calm at every $x\in\Omega$ and therefore continuous. Let :=${\cal I}:=\{t\in\mathbb{R}\,\big{|}\,x+t(x^{\prime}-x)\in\Omega\}\supset[0,1]$ and consider the function $\varphi:{\cal I}\to\mathbb{R}$, $\varphi(t):=f(x+t(x^{\prime}-x))-f(x)-t(f(x^{\prime})-f(x))$. Owing to Lemmata 3.4, 3.5, $\varphi$ is semismooth on ${\cal I}$ with respect to the mapping $\Phi(t):=\Psi(x+t(x^{\prime}-x))(x^{\prime}-x)-(f(x^{\prime})-f(x))$, $t\in{\cal I}$. There holds $\varphi(0)=\varphi(1)=0$ and, since $\varphi$ is continuous on $[0,1]$, there exists $\bar{t}\in(0,1)$ being either a global minimizer or a global maximizer of $\varphi$ on $[0,1]$. We may assume that $\bar{t}$ is a local minimizer, since otherwise we can interchange $x$ and $x^{\prime}$ to obtain a global minimizer $(1-\bar{t})$ for the function $-\varphi(1-t)$. Then
$0\in\widehat{\partial}\varphi(\bar{t})\subset(\rm cocl\,\Phi)(\bar{t})=(\rm cocl\,\Psi)(x+\bar{t}(x^{\prime}-x))(x^{\prime}-x)-(f(x^{\prime})-f(x))$
yielding the assertion.
ad(iii): This is an immediate consequence of (ii).
ad(iv): Since $f$ is locally Lipschitz on $\Omega$, for every $x\in\Omega$ there holds $\partial f(x)=(\mathrm{cl}\,\widehat{\partial}f)(x)$ by definition of $\partial f$ and $\partial_{c}f(x)=\mathrm{conv}\,\partial f(x)$ by [28, Theorem 9.61]. This yields, together with (i),
$\partial f(x)\subset\partial_{c}f(x)=(\rm cocl\,\widehat{\partial}f)(x)\subset\{g^{T}\,\big{|}\,g\in(\rm cocl\,(\rm cocl\,\Psi))(x)\}=\{g^{T}\,\big{|}\,g\in(\rm cocl\,\Psi)(x)\}.$
ad (v): In order to show that the set $\Omega\setminus\Omega_{\Psi}$ is negligible, one can take over the proof of [20, Theorem 1.12] to show that for every compact cube $K(x,\delta):=\{x^{\prime}\in\mathbb{R}^{n}\,\big{|}\,\max_{i}|x^{\prime}_{i}-x_{i}|\leq\delta\}\subset\Omega$ the set $K(x,\delta)\setminus\Omega_{\Psi}$ is negligible. Consider the countable collection $K(x_{1},\delta_{1}),K(x_{2},\delta_{2}),\ldots$ of cubes contained in $\Omega$ with rational $x_{i}$ and rational $\delta_{i}$. This collection covers $\Omega$ and therefore the set $\Omega\setminus\Omega_{\Psi}$ is negligible as the countable union of the negligible sets $K(x_{i},\delta_{i})\setminus\Omega_{\Psi}$.
For every $x\in\Omega_{\Psi}$ we conclude from (iv) that $\partial_{c}f(x)$ is a singleton and thus $f$ is strictly differentiable at $x$ with $\nabla f(x)=\{\partial_{c}f(x)\}$ by [5, Proposition 2.2.4]. Since strict differentiability of $f$ is equivalent with continuity of $\nabla f$ relative to ${\cal O}_{f}$, cf. [28, Exercise 9.64], $\nabla f$ is continuous on $\Omega_{\Psi}$. ∎
It is not difficult to extend these results to vector-valued mappings.
Proposition 3.9.
@@
    Let $\Omega_{\Psi}$ denote the set of all $x\in\Omega$ such that $(\rm cocl\,\Psi)(x)$ is a singleton. Then $\Omega\setminus\Omega_{\Psi}$ has Lebesgue measure zero. Moreover, the mapping $F$ is strictly differentiable at every $x\in\Omega_{\Psi}$ with $\{\nabla F(x)\}=(\rm cocl\,\Psi)(x)$ and the mapping $x\mapsto\nabla F(x)$ is continuous on $\Omega_{\Psi}$.
@@
ad (i): By Lemma 3.4 we obtain that each component function $F_{i}$, $i=1,\ldots,m$, is semismooth on $\Omega$ with respect to the mapping $\Psi_{i}:\Omega\rightrightarrows\mathbb{R}^{1\times n}$ given by $\Psi_{i}(x):=\{e_{i}^{T}A\,\big{|}\,A\in\Psi(x)\}$. Hence, the components $F_{i}$, $i=1,\ldots,m$ are locally Lipschitz on $\Omega$ and so is the mapping $F$ as well.
@@
ad (ii): Let $\Pi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ denote the mapping which assigns to each $x$ the set of all $m\times n$ matrices whose $i$-th row belongs to $(\rm cocl\,\Psi_{i})(x)$. Clearly, $(\rm cocl\,\Psi)(x)\subset\Pi(x)$, $x\in\mathbb{R}^{n}$. Next let $\tilde{\Omega}:=\bigcap_{i=i}^{m}\Omega_{\Psi_{i}}$. For every $x\in\tilde{\Omega}$ we obtain from Proposition 3.8(v) that for every $i=1,\ldots,m$ the component function $F_{i}$ is strictly differentiable at $x$ and $(\rm cocl\,\Psi_{i})(x)=\{\nabla F_{i}(x)\}$ is a singleton. Hence, $F$ is also strictly differentiable at $x$ and $(\rm cocl\,\Psi)(x)=\Pi(x)=\{\nabla F(x)\}$. Since the set $\Omega\setminus\tilde{\Omega}=\bigcup_{i=1}^{m}(\Omega\setminus\Omega_{\Psi_{i}})$ is negligible, the set $\tilde{\Omega}$ differs from the set ${\cal O}_{F}$, the set of all points where $F$ is differentiable, by a negligible set. By denoting
$\nabla_{s}F(x):=\begin{cases}\nabla F(x)&\mbox{if $x\in\tilde{\Omega}$,}\\ \emptyset&\mbox{else}\end{cases}$
and taking into account $\mathrm{conv}\,\overline{\nabla}F=\rm cocl\,\nabla F$, it follows from [32, Theorem 4] that $\mathrm{conv}\,\overline{\nabla}F=\rm cocl\,\nabla_{s}F$. Since $\nabla_{s}F(x)\subset(\rm cocl\,\Psi)(x)$ $\forall x\in\Omega$, we obtain $\mathrm{conv}\,\overline{\nabla}F=(\rm cocl\,\nabla_{s}F)(x)\subset\big{(}\rm cocl\,(\rm cocl\,\Psi)\big{)}(x)=(\rm cocl\,\Psi)(x)$ $\forall x\in\Omega$.

### v2

@@ 3 On $\Psi$-semismooth single-valued mappings
The following definition is closely related to the definition of semismooth operators in Banach spaces, cf. [38, Definition 3.1].
@@
    We say that $F$ is $\Psi$-ss on a subset $U\subset\Omega$ if $F$ is $\Psi$-ss at every $x\in U$. The mapping $\Psi$ is called a semismooth derivative of $F$ on $U$.
By this definition, a locally Lipschitz mapping $F$ is $\mathrm{conv}\,\overline{\nabla}F$-ss if and only if it is G-semismooth.
Our definition of semismoothness differs from Ulbrich’s one [38, Definition 3.1] insofar as we do not suppose local continuity of $F$ but require local boundedness of $\Psi$. If $F$ is $\Psi$-ss at $\bar{x}$ then $\Psi$ is also called a Newton map for $F$ at $\bar{x}$ in [20]. If $\Psi$ is single-valued then it is also called a Newton derivative (or slanting function) for $F$ at $\bar{x}$ in the literature. We dispense with using this terminology because they are normally used at single points, namely solutions of the nonsmooth equation $F(x)=0$. However, for our purposes it is essential that the properties hold not only at one single point $\bar{x}$ but also on a neighborhood of $\bar{x}$.
In the literature one can find also several papers about the $\Psi$-ss property under some additional assumptions on $\Psi$: For instance, if $F$ is $\Psi$-ss at $\bar{x}$, $\Psi(x)$ is compact for all $x$ close to $\bar{x}$ and $\Psi$ is usc at $\bar{x}$, then $F$ is called C-differentiable at $\bar{x}$ and $\Psi$ is called a C-differential operator of $F$ [33].
If a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, is $\Psi$-ss at $\bar{x}\in\Omega$, it follows immediately from the definition that each component $F_{i}$, $i=1,\ldots,m$, is semismooth at $\bar{x}$ with respect to
$\Psi_{i}(x):=\{e_{i}^{T}A\,\big{|}\,A\in\Psi(x)\},$ (3.9)
where $e_{i}$ denotes the $i$-th unit vector. Conversely. if each component $F_{i}$, $i=1,\ldots,m$, is semismooth at $\bar{x}\in\Omega$ with respect to some mapping $\Psi_{i}:\mathbb{R}^{n}\to\mathbb{R}^{1\times n}$, then $F$ is semismooth at $\bar{x}$ with respect to
$\Psi(x):=\{A\in\mathbb{R}^{m\times n}\,\big{|}\,e_{i}^{T}A\in\Psi_{i}(x)\}.$ (3.10)
This construction resembles the one of Qi’s C-subdifferential, see, e.g. [38, Proposition 2.1].
Continuously differentiable mappings $F$ are $\nabla F$-ss. Further, it is easy to see that a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, which is $\Psi$-ss at $\bar{x}\in\Omega$, is calm at $\bar{x}$ and $\mathrm{clm}\,F(\bar{x})\leq\mathrm{bnd}\,\Psi(\bar{x})$.
It is well-known that Newton maps satisfy a chain rule, see, e.g., [20, Theorem 6.14]. For the sake of completeness we state the following result, where in contrast to [20] we do not assume Lipschitz continuity of the involved mappings $F_{1},F_{2}$:
@@
By calmness of $F_{1}$ at $\bar{x}$ and calmness of $F_{2}$ at $F_{1}(\bar{x})$ we can find neighborhoods $\tilde{U}\subset U$ of $\bar{x}$ and $\tilde{V}\subset V$ of $F_{1}(\bar{x})$ together with two positive reals $\kappa_{1},\ \kappa_{2}$ such that $\|F_{1}(x)-F_{1}(\bar{x})\|\leq\kappa_{1}\|x-\bar{x}\|$ $\forall x\in\tilde{U}$, $\|F_{2}(y)-F_{2}(F_{1}(\bar{x}))\|\leq\kappa_{2}\|y-F_{1}(\bar{x})\|$ $\forall y\in\tilde{V}$ and $F_{1}(\tilde{U})\subset\tilde{V}$. By possibly shrinking $\tilde{U}$, $\tilde{V}$ and enlarging $\kappa_{1},\kappa_{2}$ we may also assume that $\mathrm{bnd}\,\Psi_{1}(x)\leq\kappa_{1}$ $\forall x\in\tilde{U}$ and $\mathrm{bnd}\,\Psi_{2}(y)\leq\kappa_{2}$ $\forall y\in\tilde{V}$. Next, for arbitrary $\epsilon>0$ we choose a neighborhood $V_{\epsilon}\subset\tilde{V}$ of $F_{1}(\bar{x})$ such that for each $y\in V_{\epsilon}$ and each $B\in\Psi_{2}(y)$ one has $\|F_{2}(y)-F_{2}(F_{1}(\bar{x}))-B(y-F_{1}(\bar{x}))\|\leq\epsilon\|y-F_{1}(\bar{x})\|$. Finally we choose a neighborhood $U_{\epsilon}\subset\tilde{U}$ such that $F_{1}(U_{\epsilon})\subset V_{\epsilon}$ and for every $x\in U_{\epsilon}$ and every $A\in\Psi_{1}(x)$ one has $\|F_{1}(x)-F_{1}(\bar{x})-A(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|$. Thus, for every $x\in U_{\epsilon}$, every $A\in\Psi_{1}(x)$ and every $B\in\Psi_{2}(F_{1}(x))$ we obtain
@@
Consider an open set $\Omega\subset\mathbb{R}^{n}$, two mappings $F_{i}:\Omega\to\mathbb{R}^{m}$ and two mappings $\Psi_{i}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m}$, $i=1,2$. If $F_{i}$ is $\Psi_{i}$-ss, $i=1,2$, at $\bar{x}\in\Omega$ then $F_{1}+F_{2}$ is $(\Psi_{1}+\Psi_{2})$-ss at $\bar{x}$.
@@
Let $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, be continuous and assume that $F$ is semismooth at $\bar{x}\in\Omega$ with respect to $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$. Then $F$ is also $\mathrm{cl}\,\Psi$-ss and ${\rm cocl}\,\Psi$-ss at $\bar{x}$.
@@
Let $\epsilon>0$ be arbitrarily fixed. By definition we can find some $\delta>0$ such that there holds
$\|F(x)-F(\bar{x})-A(x-\bar{x})\|\leq\epsilon\|x-\bar{x}\|\ \forall x\in{\cal B}_{\delta}(\bar{x})\ \forall A\in\Psi(x).$
For every $x\in{\cal B}_{\delta/2}(\bar{x})\setminus\{\bar{x}\}$ and $A\in(\mathrm{cl}\,\Psi)(x)$ we can find some sequence $(x_{k},A_{k}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,\Psi}}(x,A)$. Since $x_{k}\in{\cal B}_{\delta}(\bar{x})$ for all $k$ sufficiently large, we can conclude
$\frac{\|F(x)-F(\bar{x})-A(x-\bar{x})\|}{\|x-\bar{x}\|}=\lim_{k\to\infty}\frac{\|F(x_{k})-F(\bar{x})-A_{k}(x_{k}-\bar{x})\|}{\|x_{k}-\bar{x}\|}<\epsilon,$
verifying that $F$ is $\mathrm{cl}\,\Psi$-ss at $\bar{x}$. The ${\rm cocl}\,\Psi$-ss property of $F$ at $\bar{x}$ follows now easily from the fact that any matrix $A\in({\rm cocl}\,\Psi)(x)$ can be written as convex combination $A=\sum_{i=1}^{N}\lambda_{i}A_{i}$ with $A_{i}\in(\mathrm{cl}\,\Psi)(x)$ and $\sum_{i=1}^{N}\lambda_{i}=1$, $\lambda_{i}\geq 0$, $i=1,\ldots,N$. ∎
Remark 3.5.
As a consequence of this lemma we obtain that a continuous mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, which is $\Psi$-ss at $\bar{x}\in\Omega$, is also C-differentiable at $\bar{x}$ with C-differential operator $\mathrm{cl}\,\Psi$ or ${\rm cocl}\,\Psi$. Further, for any sequence $x_{k}\mathop{\rightarrow}\limits^{\Omega}\bar{x}$ we can select a bounded sequence $A_{k}\in\Psi(x_{k})$ and, by passing to a subsequence $x_{k_{j}}$ such that $A_{k_{j}}$ converges to some $A\in(\mathrm{cl}\,\Psi)(\bar{x})$, we obtain that
$F(x_{k_{j}})-F(\bar{x})-A(x_{k_{j}}-\bar{x})=o(\|x_{k_{j}}-\bar{x}\|),$
i.e., $F$ is H-differentiable at $\bar{x}$ with H-differential $(\mathrm{cl}\,\Psi)(\bar{x})$ in the sense of [17].
@@
The elements of $G(x)$ are called pseudogradients.
@@
Note that pseudogradients are column vectors whereas the elements of $\Psi$ are $1\times n$ matrices, i.e., row vectors.
@@
Let $\Omega\subset\mathbb{R}^{n}$ be open and let $f:\Omega\to\mathbb{R}$ be semismooth on $\Omega$ with respect to the mapping $\Psi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{1\times n}$. Then the following statements hold true.
@@
    $f$ is locally Lipschitz on $\Omega$ and for every $x\in\Omega$ there holds $\partial^{c}f(x)^{T}\subset({\rm cocl}\,\Psi)(x)$.
@@
    Let $\Omega_{\Psi}$ denote the set of all $x\in\Omega$ such that $({\rm cocl}\,\Psi)(x)$ is a singleton. Then $\Omega\setminus\Omega_{\Psi}$ has Lebesgue measure zero. Moreover, the function $f$ is strictly differentiable at every $x\in\Omega_{\Psi}$ with $\{\nabla f(x)\}=({\rm cocl}\,\Psi)(x)=\Psi(x)$.
@@
ad (i): The assertions have been shown in [27, Theorems 1.1, 1.10] for the special case $\Omega=\mathbb{R}^{n}$ and it is an easy task to carry over the proofs to general open sets $\Omega$.
ad (ii): In order to show that the set $\Omega\setminus\Omega_{\Psi}$ is negligible, one can take over the proof of [27, Theorem 1.12] to show that for every compact cube $K(x,\delta):=\{x^{\prime}\in\mathbb{R}^{n}\,\big{|}\,\max_{i}|x^{\prime}_{i}-x_{i}|\leq\delta\}\subset\Omega$ the set $K(x,\delta)\setminus\Omega_{\Psi}$ is negligible. Consider the countable collection $K(x_{1},\delta_{1}),K(x_{2},\delta_{2}),\ldots$ of cubes contained in $\Omega$ with rational $x_{i}$ and rational $\delta_{i}$. This collection covers $\Omega$ and therefore the set $\Omega\setminus\Omega_{\Psi}$ is negligible as the countable union of the negligible sets $K(x_{i},\delta_{i})\setminus\Omega_{\Psi}$.
For every $x\in\Omega_{\Psi}$ we conclude from (i) that $\partial^{c}f(x)=({\rm cocl}\,\Psi)(x)=\Psi(x)$ is a singleton and thus $f$ is strictly differentiable at $x$ with $\partial^{c}f(x)=\{\nabla f(x)\}$ by [6, Proposition 2.2.4]. ∎
We are now in the position to state the main result of this section.
Theorem 3.8.
@@
    Let $\Omega_{\Psi}$ denote the set of all $x\in\Omega$ such that $({\rm cocl}\,\Psi)(x)$ is a singleton. Then $\Omega\setminus\Omega_{\Psi}$ has Lebesgue measure zero. Moreover, the mapping $F$ is strictly differentiable at every $x\in\Omega_{\Psi}$ with $\mathrm{conv}\,\overline{\nabla}F(x)=\{\nabla F(x)\}=({\rm cocl}\,\Psi)(x)=\Psi(x)$.
@@
ad (i): Each component $F_{i}$, $i=1,\ldots,m$, is semismooth on $\Omega$ with respect to $\Psi_{i}$ given by (3.9) and therefore locally Lipschitz on $\Omega$ by Proposition 3.7. Hence, $F$ is locally Lipschitz on $\Omega$ as well.
@@
ad (ii): Let $\Pi:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{m\times n}$ denote the mapping which assigns to each $x$ the set of all $m\times n$ matrices whose $i$-th row belongs to $({\rm cocl}\,\Psi_{i})(x)$. Clearly, $({\rm cocl}\,\Psi)(x)\subset\Pi(x)$, $x\in\mathbb{R}^{n}$. Next let $\tilde{\Omega}:=\bigcap_{i=i}^{m}\Omega_{\Psi_{i}}$. For every $x\in\tilde{\Omega}$ we obtain from Proposition 3.7(ii) that for every $i=1,\ldots,m$ the component function $F_{i}$ is strictly differentiable at $x$ and $({\rm cocl}\,\Psi_{i})(x)=\{\nabla F_{i}(x)\}$ is a singleton. Hence, $F$ is also strictly differentiable at $x$ and $({\rm cocl}\,\Psi)(x)=\Pi(x)=\{\nabla F(x)\}$. Since the set $\Omega\setminus\tilde{\Omega}=\bigcup_{i=1}^{m}(\Omega\setminus\Omega_{\Psi_{i}})$ is negligible, the set $\tilde{\Omega}$ differs from the set ${\cal O}_{F}$, the set of all points where $F$ is differentiable, by a negligible set. By defining the mapping $\nabla_{s}F(x):=\nabla F(x)$ for $x\in\mathrm{dom}\,\nabla_{s}F:=\tilde{\Omega}$ and taking into account $\mathrm{conv}\,\overline{\nabla}F={\rm cocl}\,\nabla F$, it follows from [39, Theorem 4] that $\mathrm{conv}\,\overline{\nabla}F={\rm cocl}\,\nabla_{s}F$. Since $\nabla_{s}F(x)\subset({\rm cocl}\,\Psi)(x)$ $\forall x\in\Omega$, we obtain $\mathrm{conv}\,\overline{\nabla}F\subset\big{(}{\rm cocl}\,({\rm cocl}\,\Psi)\big{)}(x)=({\rm cocl}\,\Psi)(x)$ $\forall x\in\Omega$.